### v1

@@ 3.1 The construction of step, constant, and linear function approximators
We start with the equidistant case and consider four pairs of basis functions $\{\phi_{i}^{\pm}\}_{i=1}^{4}$ with locations $\{b_{i}\}_{i=1}^{4}$ and corresponding coefficients $\{p_{i},q_{i}\}_{i=1}^{4}=\{\pm b_{i}\}_{i=1}^{4}$. The locations $b_{1}\leq b_{2}\leq b_{3}\leq b_{4}$ satisfy the following symmetric distance relations,
(4) $d=b_{2}-b_{1}=b_{4}-b_{3}\,\Longrightarrow\,b_{1}+b_{4}=b_{2}+b_{3},$
where $d$ is the basis distance between the two pairs of basis functions. The step and constant function approximators share the same following form,
(5) $f_{\mu}^{\text{NN}}(x)=\sum_{i=1}^{4}p_{i}\phi_{i}^{+}(x)+\sum_{i=1}^{4}q_{i}\phi_{i}^{-}(x),\quad x\in[0,1],\quad\mu\in\{s,c\}.$
@@ 3.1.1 Step-matching construction of step function approximators $f_{s}^{\text{NN}}$
To ensure a local error of the approximator, we appeal $f_{s}^{\text{NN}}$ to be $x$-independent outside the interval $[b_{1},b_{4}]$. As a result, the coefficients $p_{i},q_{i}$ must satisfy $\sum_{i=1}^{4}p_{i}=\sum_{i=1}^{4}q_{i}=0$, which implies the correspondence between $\{p_{i},q_{i}\}_{i=1}^{4}$ and {±$\{\pm b_{i}\}_{i=1}^{4}$ as
(6) $\begin{split}p_{1}=-b_{1},\quad p_{2}=+b_{2},\quad p_{3}=+b_{3},\quad p_{4}=-b_{4},\\ q_{1}=+b_{4},\quad q_{2}=-b_{3},\quad q_{3}=-b_{2},\quad q_{4}=+b_{1}.\end{split}$
@@
(7) $\begin{split}f_{s}^{\text{NN}}(x)=\left\{\begin{array}[]{lr}2b_{1}b_{4}-2b_{2}b_{3}&0\leq x<b_{1},\vspace{0.4ex}\\ (-b_{1}+b_{4})x+b_{1}^{2}+b_{1}b_{4}-2b_{2}b_{3}&b_{1}\leq x<b_{2},\vspace{0.4ex}\\ (-2b_{1}+2b_{2})x+b_{1}^{2}-b_{2}^{2}+b_{1}b_{4}-b_{2}b_{3}&b_{2}\leq x<b_{3},\vspace{0.4ex}\\ (-b_{1}+b_{4})x+b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{1}b_{4}&b_{3}\leq x<b_{4},\vspace{0.4ex}\\ b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2}&b_{4}\leq x\leq 1.\text{ }\end{array}\right.\end{split}$
@@
(8) $h=2(b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2})=4d(b_{4}-b_{2}).$
In practice, choosing the basis functions adjacently leads to $d=\frac{1}{n-1}$ and $h=8d^{2}$. By shifting $h/2$ and scaling $\frac{\Delta h}{h}$, we use $f_{s}^{\text{NN}}$ to approach step function $h\,\chi(x-s)$ with $s\in[b_{1},b_{4}]$. It is obvious that the $L^{\infty}$ error has the following trivial bound,
(9) $\left|\left[f_{s}^{\text{NN}}(x)+\frac{h}{2}\right]-h\,\chi(x-s)\right|\leq h,\quad\forall\,x\in[0,1].$
Eq. (7) implies that the approximation is exactly accurate when $x\notin[b_{1},b_{4}]$.
Remark 3.1.
Although it was suggested that a step function could be approached by two ReLU basis functions [5], this approach is unsuitable for the permutation training scenario. For an error tolerance $\varepsilon$, a step function $f_{s}(x)=h_{s}\chi(x-s)$ can be well approximated by a linear combination of two ReLU basis functions as
@@ 3.1.2 Constant-matching construction of constant function approximators $f_{c}^{\text{NN}}$
The construction of $f_{c}^{\text{NN}}$ shares the same form of $f_{s}^{\text{NN}}$ in Eq. (5) but differs slightly in the parameter assignment in Eq. (6). The coefficients $\{p_{i},q_{i}\}_{i=1}^{4}$ are set to equalize the height of the two constant pieces $x<b_{1}$ and $b_{4}\leq x$, leading to $-\sum_{i=1}^{4}p_{i}b_{i}=\sum_{i=1}^{4}q_{i}b_{i}$. A possible choice is
(10) $\begin{split}p_{1}=-b_{1},\quad p_{2}=+b_{2},\quad p_{3}=+b_{3},\quad p_{4}=-b_{4},\\ q_{1}=+b_{1},\quad q_{2}=-b_{2},\quad q_{3}=-b_{3},\quad q_{4}=+b_{4}.\end{split}$
@@
(11) $\begin{split}f_{c}^{\text{NN}}(x)&=(b_{2}+b_{3}-b_{1}-b_{4})\,x+b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2}\\ &=2d(b_{4}-b_{2})=h/2,\quad x\in[0,1].\end{split}$
@@
Therefore, we can freely adjust the sign $a_{j}=\pm 1$ of the approached constant. It also enables us to construct $f_{c}^{\text{NN}}$ and $f_{-c}^{\text{NN}}$ with the same constant separately, then pair them up to offset with each other, i.e., $f_{c}^{\text{NN}}(x)+f_{-c}^{\text{NN}}(x)=0$ for all $x\in[0,1]$.
@@ 3.1.3 Linear reorganization of the linear function approximators $f_{\ell}^{\text{NN}}$
(12) $f_{\ell}^{\text{NN}}(x)=m_{i}\ell_{i},\quad\ell_{i}(x):=b_{i}\phi_{i}^{+}(x)-b_{i}\phi_{i}^{-}(x)=b_{i}x-b_{i}^{2},\quad x\in[0,1],$
@@ 3.2 Annihilate the unused part of the network
(13) $\mathcal{S_{\ell}}(x)=\sum_{i\in I_{\text{un}}}f_{\ell_{i}}^{\text{NN}}(x)=\sum_{i\in I_{\text{un}}}m_{i}b_{i}x-\sum_{i\in I_{\text{un}}}m_{i}b_{i}^{2}=:\beta x+\eta,\quad x\in[0,1],$
where $\beta$ is the slope and $\eta$ is the intercept. The goal then is to control the $L^{\infty}$-norm of $\mathcal{S_{\ell}}$. We first choose a proper $m_{i}\in\{-1,1\}$ for each $i\in I_{\text{un}}$ to reduce the slope $\beta$, which is equivalent to assigning addition and subtraction operations within a given sequence to reduce the final result. Motivated by the Leibniz’s test (known as alternating series test) [36], the following lemma offers a solution with an upper bound of $\beta$.
@@
(14) $0\leq\sum_{i=1}^{\bar{n}}m_{i}c_{i}\leq\Delta c,\quad\Delta c=\max_{i}\left(\min_{j\neq i}|c_{i}-c_{j}|\right),$
where $\Delta c$ can be regarded as the largest gap between the adjacent elements in $\{c_{i}\}_{i=1}^{\bar{n}}$.
Proof 3.3.
Without loss of generality, we assume $c_{1}\geq c_{2}\geq\cdots\geq c_{\bar{n}}\geq 0$ and thus $\Delta c=\max_{i}|c_{i}-c_{i-1}|$. Since $\bar{n}$ is an even number, we note $\bar{n}=2\tilde{n}$ and define a new series $\{r_{j}\}_{j=1}^{\tilde{n}}$ as the difference between each pair of elements in $\{c_{i}\}_{i=1}^{\bar{n}}$,
$r_{j}=c_{2j-1}-c_{2j}\geq 0,\quad j=1,2,\cdots,\tilde{n}.$
Then we permute $(r_{j})_{j=1}^{\tilde{n}}$ to be descending order. Concretely, we find a permutation $\tau$ and denote $r^{\prime}_{j}=r_{\tau(j)}=(c_{\tau(2j-1)}-c_{\tau(2j)})$, such that $r^{\prime}_{1}\geq r^{\prime}_{2}\geq\cdots\geq r^{\prime}_{\tilde{n}}\geq 0$. Next, we alternatively use addition and subtraction operations on each $r_{j}^{\prime}$ by noting $\lambda_{j}=(-1)^{j-1}$, $j=1,\cdots,\tilde{n}$, and estimate the summation $\mathcal{S}_{\tilde{n}}$ as the following forms,
$\mathcal{S}_{\tilde{n}}:=\sum_{j=1}^{\tilde{n}}\lambda_{j}r^{\prime}_{j}=\left\{\begin{array}[]{l}(r_{1}^{\prime}-r_{2}^{\prime})+(r_{3}^{\prime}-r_{4}^{\prime})+\cdots\geq 0,\\ r_{1}^{\prime}-(r_{2}^{\prime}-r_{3}^{\prime})-(r_{4}^{\prime}-r_{5}^{\prime})-\cdots\leq r_{1}^{\prime}\leq\Delta c.\end{array}\right.$
Note that $\mathcal{S}_{\tilde{n}}=\sum_{j=1}^{\tilde{n}}\lambda_{j}(c_{\tau(2j-1)}-c_{\tau(2j)})$ is of the form of $\sum_{i=1}^{\bar{n}}m_{i}c_{i}$, hence the choice of $m_{i}$ implied by $\lambda_{j}$ satisfies our requirement and the proof is finished.
Remark 3.4.
@@
Proof 3.5 (Proof of Theorem 2.2).
For any given target function $f^{*}\in C([0,1])$ and small number $\varepsilon$, we aim to construct a network $f^{\text{NN}}$ with $n$ pairs of basis functions to approximate $f^{*}$ with error $\varepsilon$. The goal is achieved with the following points:
a) Approach $f^{*}$ by a piecewise constant function $g$. Employing Lemma 2.5, we can construct a piecewise constant function $g$ in Eq. (3) with a constant height $\Delta h\leq\varepsilon/4$, distinct step locations $s_{1}<\cdots<s_{J}$, and $\delta_{s}=\max_{j}|s_{j}-s_{j-1}|$ as the maximal gap of step locations. It gives $|g(x)-f^{*}(x)|\leq\Delta h<\varepsilon/2$ for all $x\in[0,1]$.
@@
However, to maintain the consistency of the following estimation, we refine the basis locations $B^{(\hat{n})}$ to $B^{(n)}$ with $n=L(\hat{n}-1)+1$ for some integer $L\geq\Delta h(\hat{n}-1)^{2}/8$ (see Fig. 1(c)). Denote $K=\{1,\cdots,n\}$ as the index set of the locations $B^{(n)}$, and for each $j=1,\cdots,J$, we choose the basis functions with the index $K_{j}:=\{k_{j},k_{j}+L,k_{j}+2L,k_{j}+3L\}\subset K$, such that $s_{j}\in[b_{k_{j}+L},b_{k_{j}+2L})$ and $\{K_{j}\}_{j=1}^{J}$ has empty intersection. As a consequence, $\{p_{k},q_{k}\}_{k\in K_{j}}$ and {±$\{\pm b_{k}\}_{k\in K_{j}}$ are assigned to be step-matching (according to Eq. (6)) to construct $\frac{\Delta h}{h}(f_{s_{j}}^{\text{NN}}+a_{j}\frac{h}{2})$ as a subnetwork of $f^{\text{NN}}$ to approximate $a_{j}f_{s_{j}}$ with the following error estimation,
(15) $\left|\frac{\Delta h}{h}\left[f_{s_{j}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}f_{s_{j}}(x)\right|\leq\Delta h,\quad\forall\,x\in[0,1],\quad j=1,\cdots,J,$
where $h=8\hat{d}^{2}=8/(\hat{n}-1)^{2}=8L^{2}/(n-1)^{2}$ is the height determined by $n$, and the scaling $\Delta h/h=:\gamma$ serves to match the constant height $\Delta h$ and its approximator height $h$. Our requirement of $L$ gives the condition $L\geq\gamma$. This estimation is following Eq. (9) with additional $a_{j}$. Furthermore, denote $K_{\text{use}}=\cup_{j=1}^{J}K_{j}$ as the index set for all involved basis functions in Eq. (15), which leads to the requirement of $\hat{n}>4J$. We define $\mathcal{S}_{\text{use}}$ as the picked subnetwork of $f^{\text{NN}}$,
@@
(16) $\begin{split}E_{\text{use}}:=&\,\max_{x\in[0,1]}\left|\mathcal{S}_{\text{use}}(x)+\frac{h}{2}\sum_{j=1}^{J}a_{j}-\frac{h}{\Delta h}g(x)\right|\\ =&\,\max_{x\in[0,1]}\left|\sum_{j=1}^{J}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-\sum_{j=1}^{J}\frac{h}{\Delta h}a_{j}f_{s_{j}}(x)\right|\leq h.\end{split}$
@@
$\mathcal{S_{\text{un}}}(x)=\sum_{k\in K\setminus K_{\text{use}}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{k\in K\setminus K_{\text{use}}}(m_{k}b_{k}x-m_{k}b_{k}^{2})=:\beta x+\eta.$
Since the number of unused basis functions can be processed to be even, we apply Lemma 3.2 on the series $\{b_{k}\}_{k\in K\setminus K_{\text{use}}}$ to get a choice of $m_{k}\in\{-1,1\}$ for each $k\in K\setminus K_{\text{use}}$, which provides an upper bound of the slope $\beta$ as $0\leq\beta\leq\Delta b$, where $\Delta b$ is the largest gap between the adjacent locations in $\{b_{k}\}_{k\in K\setminus K_{\text{use}}}$. Notice that in $f^{\text{NN}}$, the refined basis distance $d=\frac{1}{n-1}<\frac{1}{\gamma\hat{n}}$ is small enough to ensure that there is at least one unused basis function between two adjacent used basis functions, i.e., $\Delta b\leq\frac{2}{\gamma\hat{n}}$. To control the error of the intercept $\eta$, we introduce a shifting $C_{\eta}=-\eta$, thus the error $E_{\text{un}}$ introduced by the unused part gives the following estimation,
(17) $E_{\text{un}}:=\max_{x\in[0,1]}\big{|}\mathcal{S_{\text{un}}}(x)+C_{\eta}\big{|}\leq\Delta b<\frac{2}{\gamma\hat{n}}=\frac{2h}{\Delta h\hat{n}}.$
@@
This complete the proof since $|f^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.

### v2

@@ 3.1 The construction of step, constant, and linear function approximators
Lemma 2.1 enables us to approximate the target function $f^{*}$ by a piecewise constant function $g$ in Eq. (3). Next, we aim to approximate each step function $f_{s_{j}}$ within $g$, respectively, and then eliminate the remaining part of the network. This section introduces several key constructions to achieve this goal in the equidistant case.
@@ 3.1.1 Step-matching construction of step function approximators $f_{s}^{\text{NN}}$
Here we construct the step function approximator $f_{s}^{\text{NN}}$ for a given step function $f_{s}(x)=\Delta h\,\chi(x-s)$ with height $\Delta h$ and location $s$. The construction considers four pairs of basis functions $\{\phi_{i}^{\pm}\}_{i=1}^{4}$ with locations $\{b_{i}\}_{i=1}^{4}$ and coefficients $\{p_{i},q_{i}\}_{i=1}^{4}=\{\pm b_{i}\}_{i=1}^{4}$, which has the following form,
$f_{s}^{\text{NN}}(x)=\sum_{i=1}^{4}p_{i}\phi_{i}^{+}(x)+\sum_{i=1}^{4}q_{i}\phi_{i}^{-}(x),\quad x\in[0,1].$ (4)
Here we require the locations $b_{1}<b_{2}<b_{3}<b_{4}$ satisfy the following symmetric condition,
$d:=b_{2}-b_{1}=b_{4}-b_{3}\,\Longrightarrow\,b_{1}+b_{4}=b_{2}+b_{3},$ (5)
where $d$ is the basis distance. To ensure a local error of the approximator, we appeal $f_{s}^{\text{NN}}$ to be $x$-independent outside the interval $[b_{1},b_{4}]$. As a result, the coefficients $p_{i},q_{i}$ must satisfy $\sum_{i=1}^{4}p_{i}=\sum_{i=1}^{4}q_{i}=0$, which implies the correspondence between $\{p_{i},q_{i}\}_{i=1}^{4}$ and {±$\{\pm b_{i}\}_{i=1}^{4}$ as
$\begin{split}p_{1}=-b_{1},\quad p_{2}=+b_{2},\quad p_{3}=+b_{3},\quad p_{4}=-b_{4},\\ q_{1}=+b_{4},\quad q_{2}=-b_{3},\quad q_{3}=-b_{2},\quad q_{4}=+b_{1},\end{split}$ (6)
@@
$\begin{split}f_{s}^{\text{NN}}(x)=\left\{\begin{array}[]{lr}2b_{1}b_{4}-2b_{2}b_{3}&0\leq x<b_{1},\vspace{0.4ex}\\ (-b_{1}+b_{4})x+b_{1}^{2}+b_{1}b_{4}-2b_{2}b_{3}&b_{1}\leq x<b_{2},\vspace{0.4ex}\\ (-2b_{1}+2b_{2})x+b_{1}^{2}-b_{2}^{2}+b_{1}b_{4}-b_{2}b_{3}&b_{2}\leq x<b_{3},\vspace{0.4ex}\\ (-b_{1}+b_{4})x+b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{1}b_{4}&b_{3}\leq x<b_{4},\vspace{0.4ex}\\ b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2}&b_{4}\leq x\leq 1.\text{ }\end{array}\right.\end{split}$ (7)
@@
$h=2(b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2})=4d(b_{4}-b_{2}).$ (8)
Notice that choosing the basis functions adjacently will lead to $d=\frac{1}{n-1}$ and $h=8d^{2}$. By shifting $h/2$ and scaling $\Delta h/h$, we use $f_{s}^{\text{NN}}$ to approach step function $f_{s}$ with $s\in[b_{1},b_{4}]$. It is obvious that the $L^{\infty}$ error has the following trivial bound,
$\left|\frac{\Delta h}{h}\left[f_{s}^{\text{NN}}(x)+\frac{h}{2}\right]-f_{s}(x)\right|\leq\Delta h,\quad\forall\,x\in[0,1].$ (9)
Eq. (7) implies that the approximation is exactly accurate when $x\notin[b_{1},b_{4}]$. A toy example of this step-matching construction is shown in F.
Remark 1.
Although Cai et al. (2021) suggested that a step function could be approached by two ReLU basis functions, this approach is unsuitable for the permutation training scenario. For an error tolerance $\varepsilon$, a step function $f_{s}(x)=h_{s}\chi(x-s)$ can be well approximated by a linear combination of two ReLU basis functions as
@@ 3.1.2 Constant-matching construction of constant function approximators $f_{c}^{\text{NN}}$
The four-pair form in Eq. (4) can also be utilized to approximate a constant function, which plays a crucial role in the proof of Theorem 2.2. The constant function approximator (denoted as $f_{c}^{\text{NN}}$) shares the same form of $f_{s}^{\text{NN}}$ in Eq. (4) but differs slightly in the parameter assignment in Eq. (6). The coefficients $\{p_{i},q_{i}\}_{i=1}^{4}$ are set to equalize the height of the two constant pieces $x<b_{1}$ and $b_{4}\leq x$, leading to $-\sum_{i=1}^{4}p_{i}b_{i}=\sum_{i=1}^{4}q_{i}b_{i}$. A possible choice is
$\begin{split}p_{1}=-b_{1},\quad p_{2}=+b_{2},\quad p_{3}=+b_{3},\quad p_{4}=-b_{4},\\ q_{1}=+b_{1},\quad q_{2}=-b_{2},\quad q_{3}=-b_{3},\quad q_{4}=+b_{4}.\end{split}$ (10)
@@
$\begin{split}f_{c}^{\text{NN}}(x)&=(b_{2}+b_{3}-b_{1}-b_{4})\,x+b_{1}^{2}-b_{2}^{2}-b_{3}^{2}+b_{4}^{2}\\ &=2d(b_{4}-b_{2})=h/2,\quad x\in[0,1].\end{split}$ (11)
@@
Therefore, we can freely adjust the sign $a_{j}=\pm 1$ of the approached constant. It also enables us to construct $f_{c}^{\text{NN}}$ and $f_{-c}^{\text{NN}}$ with the same constant separately, then pair them up to offset with each other, i.e., $f_{c}^{\text{NN}}(x)+f_{-c}^{\text{NN}}(x)=0$ for all $x\in[0,1]$. A toy example of this constant-matching construction is shown in F.
@@ 3.1.3 Linear reorganization of the linear function approximators $f_{\ell}^{\text{NN}}$
$f_{\ell}^{\text{NN}}(x)=m_{i}\ell_{i},\quad\ell_{i}(x):=b_{i}\phi_{i}^{+}(x)-b_{i}\phi_{i}^{-}(x)=b_{i}x-b_{i}^{2},\quad x\in[0,1],$ (12)
@@ 3.2 Annihilate the unused part of the network
$\mathcal{S_{\ell}}(x)=\sum_{i\in I_{\text{un}}}f_{\ell_{i}}^{\text{NN}}(x)=\sum_{i\in I_{\text{un}}}m_{i}b_{i}x-\sum_{i\in I_{\text{un}}}m_{i}b_{i}^{2}=:\beta x+\eta,\quad x\in[0,1],$ (13)
where $\beta$ is the slope and $\eta$ is the intercept. The goal then is to control the $L^{\infty}$-norm of $\mathcal{S_{\ell}}$. We first choose a proper $m_{i}\in\{-1,1\}$ for each $i\in I_{\text{un}}$ to reduce $|\beta|$, which is equivalent to assigning addition and subtraction operations within a given sequence to reduce the final result’s absolute value. Motivated by the Leibniz’s test (known as alternating series test) (Rudin, 1976), the following lemma offers a solution with an upper bound of $\beta$.
@@
$0\leq\sum_{i=1}^{\bar{n}}m_{i}c_{i}\leq\Delta c,\quad\Delta c=\max_{\begin{subarray}{c}1\leq j\leq\bar{n}\\ j\notin\operatorname{argmax}c_{j}\end{subarray}}\left(\min_{\begin{subarray}{c}i\neq j\\ c_{i}\geq c_{j}\end{subarray}}(c_{i}-c_{j})\right),$ (14)
where $\Delta c$ can be regarded as the largest gap between the adjacent elements after sorting $c_{i}$ in descending order (see G for a toy example).
Proof.
Without loss of generality, we assume $c_{1}\geq c_{2}\geq\cdots\geq c_{\bar{n}}\geq 0$ and thus $\Delta c=\max_{i}|c_{i}-c_{i-1}|$. Since $\bar{n}$ is an even number, we define a new series $\{r_{j}\}_{j=1}^{\bar{n}/2}$ as the difference between each pair of elements in $\{c_{i}\}_{i=1}^{\bar{n}}$,
$r_{j}=c_{2j-1}-c_{2j}\geq 0,\quad j=1,2,\cdots,\bar{n}/2.$
Then we permute $(r_{j})_{j=1}^{\bar{n}/2}$ to be descending order. Concretely, we find a permutation $\tau$ and denote $r^{\prime}_{j}=\tau(r_{j})=[\tau(c_{2j-1})-\tau(c_{2j})]$, such that $r^{\prime}_{1}\geq r^{\prime}_{2}\geq\cdots\geq r^{\prime}_{\bar{n}/2}\geq 0$. Next, we alternatively use addition and subtraction operations on each $r_{j}^{\prime}$ by noting $\lambda_{j}=(-1)^{j-1}$, $j=1,\cdots,\bar{n}/2$, and estimate the summation $\mathcal{S}_{\bar{n}/2}$ as the following forms,
$\mathcal{S}_{\bar{n}/2}:=\sum_{j=1}^{\bar{n}/2}\lambda_{j}r^{\prime}_{j}=\left\{\begin{array}[]{l}(r_{1}^{\prime}-r_{2}^{\prime})+(r_{3}^{\prime}-r_{4}^{\prime})+\cdots\geq 0,\\ r_{1}^{\prime}-(r_{2}^{\prime}-r_{3}^{\prime})-(r_{4}^{\prime}-r_{5}^{\prime})-\cdots\leq r_{1}^{\prime}\leq\Delta c.\end{array}\right.$
Note that $\mathcal{S}_{\bar{n}/2}=\sum_{j=1}^{\bar{n}/2}\lambda_{j}[\tau(c_{2j-1})-\tau(c_{2j})]$ is of the form of $\sum_{i=1}^{\bar{n}}m_{i}c_{i}$, hence the choice of $m_{i}$ implied by $\lambda_{j}$ satisfies our requirement and the proof is finished ∎
Remark 2.
@@
Proof of Theorem 2.1.
For any target function $f^{*}\in C([0,1])$ and small number $\varepsilon$, we aim to construct a network $f^{\text{NN}}$ with $n$ pairs of basis functions to approximate $f^{*}$ with error $\varepsilon$. The goal is achieved with the following points:
a) Approach $f^{*}$ by a piecewise constant function $g$. Employing Lemma 2.1 by letting $\varepsilon^{\prime}=\varepsilon/4$, we can construct a piecewise constant function $g$ in Eq. (3) with a constant height $\Delta h\leq\varepsilon^{\prime}=\varepsilon/4$, distinct step locations $s_{1}<\cdots<s_{J}$. It gives $|g(x)-f^{*}(x)|\leq\Delta h<\varepsilon/2$ for all $x\in[0,1]$. Here we denote $\delta_{s}:=\max_{j}|s_{j}-s_{j-1}|$ as the maximal gap of step locations, where $j=0,1,\cdots,J+1$ and $s_{0}=0$, $s_{J+1}=1$.
@@
However, to maintain the consistency of the following estimation, we refine the basis locations $B^{(\hat{n})}$ to $B^{(n)}$ with $n=L(\hat{n}-1)+1$ for some integer $L\geq\Delta h(\hat{n}-1)^{2}/8$ (see Fig. 1(c)). Denote $K=\{1,\cdots,n\}$ as the index set of the locations $B^{(n)}$, and for each $j=1,\cdots,J$, we choose the basis functions with the index $K_{j}:=\{k_{j},k_{j}+L,k_{j}+2L,k_{j}+3L\}\subset K$, such that $s_{j}\in[b_{k_{j}+L},b_{k_{j}+2L})$ and $\{K_{j}\}_{j=1}^{J}$ has empty intersection. Therefore, for each $a_{j}f_{s_{j}}$, an approximator $(f_{s_{j}}^{\text{NN}}+a_{j}h/2)$ with a shifting $a_{j}h/2$ can be constructed by applying the step-matching construction on $\{p_{k},q_{k}\}_{k\in K_{j}}$. This approximation has the following error estimation given by Eq. (9):
$\left|\left[f_{s_{j}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}\frac{h}{\Delta h}f_{s_{j}}(x)\right|\leq h,\quad\forall\,x\in[0,1],\quad j=1,\cdots,J,$ (15)
where $h=8\hat{d}^{2}=8/(\hat{n}-1)^{2}=8L^{2}/(n-1)^{2}$ is the height determined by $\hat{n}$, and the scaling $h/\Delta h$ serves to match the constant height $\Delta h$ and its approximator height $h$. Our requirement of $L$ gives the condition $L\geq\Delta h/h$. Furthermore, denote $K_{\text{use}}=\cup_{j=1}^{J}K_{j}$ as the index set for all involved basis functions in Eq. (15), which leads to the requirement of $\hat{n}>4J$. We define $\mathcal{S}_{\text{use}}$ as the picked subnetwork of $f^{\text{NN}}$,
@@
$\begin{split}E_{\text{use}}:=&\,\max_{x\in[0,1]}\left|\mathcal{S}_{\text{use}}(x)+\frac{h}{2}\sum_{j=1}^{J}a_{j}-\frac{h}{\Delta h}g(x)\right|\\ =&\,\max_{x\in[0,1]}\left|\sum_{j=1}^{J}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-\sum_{j=1}^{J}\frac{h}{\Delta h}a_{j}f_{s_{j}}(x)\right|\leq h.\end{split}$ (16)
@@
$\mathcal{S_{\text{un}}}(x)=\sum_{k\in K\setminus K_{\text{use}}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{k\in K\setminus K_{\text{use}}}(m_{k}b_{k}x-m_{k}b_{k}^{2})=:\beta x+\eta,$
Since the number of unused basis functions can be processed to be even, we apply Lemma 3.1 on the series $\{b_{k}\}_{k\in K\setminus K_{\text{use}}}$ to get a choice of $m_{k}\in\{-1,1\}$ for each $k\in K\setminus K_{\text{use}}$, which provides an upper bound of the slope $\beta$ as $0\leq\beta\leq\Delta b$, where $\Delta b$ is the largest gap between the adjacent locations in $\{b_{k}\}_{k\in K\setminus K_{\text{use}}}$. Notice that in $f^{\text{NN}}$, the refined basis distance $d<1/L\hat{n}$ is small enough to ensure that there is at least one unused basis function between two adjacent used basis functions, i.e., $\Delta b\leq 2/L\hat{n}$. To control the error of the intercept $\eta$, we introduce a shifting $C_{\eta}=-\eta$, thus the error $E_{\text{un}}$ introduced by the unused part gives the following estimation,
$E_{\text{un}}:=\max_{x\in[0,1]}\big{|}\mathcal{S_{\text{un}}}(x)+C_{\eta}\big{|}\leq\Delta b<\frac{2}{L\hat{n}}\leq\frac{2h}{\Delta h\hat{n}}.$ (17)
@@
This complete the proof since $|f^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$. ∎